# TF-Restormer: Complex Spectral Prediction for Speech Restoration

### v2

@@ 5.3 Ablation study
Effects of scaled log-spectral loss.
First, we compare the scaled log-spectral (S-log) loss against two other regression losses as a auxiliary objective to the perceptual loss: $\ell_{1}$ loss on magnitude of STFT (Mag.-only) as in (Babaev et al., 2024) and $\ell_{1}$ loss on real, imaginary, and magnitude (Cplx.) for STFT. In Table 4(a), using a complex loss shows the better results than only guiding magnitude component for all the cases. Also, using the proposed scaled log-spectral loss further improved the performance by stabilizing the gradients, confirming that finer spectral detail can be preserved without sacrificing global fidelity.
Decoder design.
Next, we validate our encoder-decoder design when $f_{E}\neq f_{D}$. Compared to the proposed method with MHCA, we test two variants: (i) a large decoder without an encoder, where extension queries are padded at the input, and (ii) an encoder-decoder without MHCA, where the decoder shares the same block design as the encoder. All three cases converge to similar behavior when $f_{E}=f_{D}$. For fairness, we rescaled the decoder-only case to match computation and memory. Illustrations are provided in Appendix E. As shown in Table 4(b), the encoder-decoder design yields more stable improvements by focusing computation on the input bandwidth. The decoder-only case produces unstable results in full-band reconstruction ($8\to 44.1$kHz). Adding the cross-self frequency module further improves performance: high-frequency bands are reconstructed more reliably via cross-attention using encoder features, whereas the isotropic encoder-decoder relies only on self-attention and thus performs worse.
Effects of frequency projection.
Finally, we examine the effect of the frequency projection module on robustness. Without projection, RoPE provides only positional context, resulting in the conventional symmetric TF dual-path structure. In contrast, adding projection to F-MHSA imposes structural bias along the frequency axis, yielding more stable training and consistent improvements across tasks, particularly under severe distortions and during high-frequency extension. These results demonstrate that structural bias is valuable for robust restoration under extreme conditions.
@@ 6 Conclusion
We presented TF-Restormer, a speech restoration model with an encoder-decoder design that enables efficient operation across arbitrary input and output sampling rates. By concentrating on the input bandwidth with a strong TF dual-path encoder and extending high frequencies through lightweight decoder queries with a cross-self mechanism, TF-Restormer achieves balanced improvements in both signal- and semantic-level fidelity while also showing robust performance on single-task benchmarks such as denoising and bandwidth extension. We also proposed a shared SFI-STFT discriminator for unified adversarial training across diverse sampling rates. Together, these results establish TF-Restormer as a universal and practical framework for robust speech restoration across diverse sampling conditions.
Reproducibility statement
We make our implementation code and demo anonymously accessible. Our anonymous link for pre-trained TF-Restormer and inference code is also included in demo pages: https://tf-restormer.github.io/demo
Ethics statement
@@
Appendix A Use of Large Language Models
This paper was written by the authors. Large language models (LLMs) were used only for minor language polishing.
Appendix B Details of Training Procedure
B.1 Simulation of training dataset
@@
B.2 Training Details
For pretraining, TF-Restormer was optimized with batch size set to 2 to train on a single GPU of NVIDIA RTX 6000 Ada Generation 48GB and AdamW (Loshchilov & Hutter, 2019) as our optimizer. For pretraining, we update 200,000 steps with VCTK dataset. The length of training utterance is set to 3 seconds. Then, adversarial training is followed with 200,000 steps. For both the stages, we use learning rate 2.0e-4 with betas (0.9, 0.995) and learning rate decay 0.9 at every 10,000 steps after 100,000 steps for pretraining whereas 10,000 steps for adversarial training.
Each stage is trained with a linear warm-up for 5,000 steps for the generator. On the other hand, the discriminators in the adversarial training perform 2 optimizations for every single generator’s optimization step without warm-up. For the discriminators, the AdamW optimizer is used with betas (0.8, 0.999). Based on conventional multi-scale STFT discriminators (Défossez et al., 2023), we used the proposed multi-scale SFI-STFT discriminator with STFT window durations of [20, 40, 60, 80, 100] ms, which allows for the capture of spectral information at different resolutions.
Appendix C Details of Model Configuration
@@
C.1 Comparison of the model size and RTF
In Table 6, we compare model size and multiply-accumulate operations (MACs) for a 1-second-long input using ptflops package [^2] https://github.com/sovrasov/flops-counter.pytorch. We also measure real-time factor (RTF) measured on 4-second-long samples with an NVIDIA RTX 4090. Conventional models operate at fixed input-output sampling rates, which results in fixed MACs regardless of the task configuration. In contrast, TF-Restormer adapts its computation depending on the input and output rates $f_{E}$ and $f_{D}$.
@@ C.1 Comparison of the model size and RTF
Appendix D Simulation of VCTK Noisy Distorted Input
The noisy-distorted input from VCTK testset in Table 3 was generated by corrupting clean VCTK utterances with additive noise from DEMAND (Thiemann et al., 2013) and colored Gaussian noise, RIR samples from RWCP (Nakamura et al., 2000) and AIR (Jeub et al., 2009) for reverberation, and distortions such as clipping and band-limiting. Additional digital effects including audio codecs (MP3, OGG) were applied before resampling to various rates (8-48 kHz). This simulation aligns the training pipeline while maintaining samples partitioning of speech, noise, and RIR sources.
@@ Appendix D Simulation of VCTK Noisy Distorted Input
Appendix E Ablation Details
E.1 Effects of scaled log-spectral loss
@@
E.2 Illustration of encoder-decoder structure ablation
@@
Appendix F Evaluation on Real-recorded Dataset
VoxCeleb Data
We test TF-Restormer on 50 selected utterances from VoxCeleb1 (Nagrani et al., 2017), balanced across gender and covering the Speech Transmission Index (STI) range 0.75-0.99, to examine performance on real-recorded speech with diverse speakers (Su et al., 2020). For VoxCeleb Data, our TF-Restormer shows as natural restoration outputs as recent competing vocoder or diffusion-based methods in terms of UTMOS, WVMOS, and DNSMOS.
Table 8: Comparison with prior work on Voxceleb real data.
VoxCeleb (HiFi-GAN-2 validation set, real data)
Model UTMOS ($\uparrow$) WVMOS ($\uparrow$) DNSMOS ($\uparrow$)
Input 2.76 $\pm$ 0.13 2.90 $\pm$ 0.16 2.72 $\pm$ 0.11
HiFi++ (Andreev et al., 2022) 2.76 $\pm$ 0.13 2.68 $\pm$ 0.14 2.98 $\pm$ 0.07
VoiceFixer 2.60 $\pm$ 0.09 2.79 $\pm$ 0.09 3.08 $\pm$ 0.06
DEMUCS (Défossez et al., 2019) 3.51 $\pm$ 0.08 3.72 $\pm$ 0.08 3.27 $\pm$ 0.04
StoRM 3.29 $\pm$ 0.08 3.54 $\pm$ 0.09 3.17 $\pm$ 0.04
BBED (Lay et al., 2023) 3.30 $\pm$ 0.10 3.47 $\pm$ 0.08 3.23 $\pm$ 0.04
HiFi-GAN-2 (Su et al., 2021) 3.67 $\pm$ 0.09 3.96 $\pm$} 0.06 3.32 $\pm$} 0.03
@@ Appendix F Evaluation on Real-recorded Dataset
TF-Restormer 3.98 $\pm$ 0.09 3.82 $\pm$ 0.09 3.34 $\pm$ 0.06
TF-Restormer-streaming 3.64 $\pm$ 0.10 3.59 $\pm$ 0.12 3.29 $\pm$ 0.04
Appendix G Limitations and Future Works
Although TF-Restormer shows balanced improvements in signal fidelity and perceptual quality, several limitations remain. First, when the input speech is severely degraded, the high uncertainty in the observations can induce content or speaker artefacts. This reflects a common trade-off in speech restoration: adversarial training encourages perceptually natural reconstructions by sampling plausible high-frequency details, but can deviate from the exact clean reference, whereas purely supervised objectives stay closer to the ground truth yet tend to average out details, producing less natural-sounding results. Second, the model may exhibit language dependency. Because it is trained on extremely distorted scenarios with limited single-language datasets, the model (and WavLM used for perceptual loss) may amplify language-specific biases (e.g. memorizing certain phonetic patterns).

### v3

@@ 5.4 Ablation study
Encoder-decoder design
In Table 5, we first analyze the contribution of the encoder–decoder structure (See Appendix E for detailed illustration.). The Encoder-only model removes the decoder entirely and applies nine encoder blocks after padding the extension queries at the input. Although this variant has a small parameter count (11.6M), its MACs are extremely large (151-415G), since the encoder must jointly infer the observed low band and synthesize the missing high-frequency components. The w/o MHCA variant restores the encoder–decoder structure but replaces the freq cross-self module with the freq. self module used in the encoder, preventing the decoder from conditioning on encoder features. Our proposed w/ MHCA model incorporates cross–self frequency attention, enabling more reliable reconstruction through encoder-conditioned queries.
To further isolate the effect of architectural design from model size, we additionally include a reduced version of the proposed model (w/ MHCA (S)), whose parameter count matches that of the encoder-only configuration. Despite having far fewer MACs than encoder-only, this size-matched variant consistently outperforms the encoder-only model across all bandwidth settings (Table 5(b)), confirming that the gains arise from the explicit separation of analysis and reconstruction and the use of cross-attention rather than increased parameter count or computational cost.
Table 6: Ablation study on SFI-discriminator and frequency projection layer in freq. module
STFT-Disc. PESQ UTMOS
UNIVERSE (GSR)
Separate 2.27 4.08
Shared (SFI) 2.29 4.10
VCTK+DEMAND (DN)
Separate 3.28 4.07
Shared (SFI) 3.41 4.14
VCTK (SSR., 8 $\to$ 16kHz)
Separate 3.64 4.10
Shared (SFI) 3.70 4.11
(a) Effect of discriminator
Case Size(M) PESQ UTMOS
UNIVERSE (GSR)
w/o F-proj. 28.1 2.26 3.90
w/ F-proj.(sep.) 63.6 2.31 4.11
w/ F-proj.(sha.) 30.1 2.29 4.10
VCTK+DEMAND (DN)
w/o F-proj. 28.1 3.21 4.03
w/ F-proj.(sep.) 63.6 3.38 4.15
w/ F-proj.(sha.) 30.1 3.41 4.14
VCTK (SSR, 8 $\to$ 16kHz)
w/o F-proj. 28.1 3.54 3.97
w/ F-proj.(sep.) 63.6 3.545 3.94
w/ F-proj.(sha.) 30.1 3.70 4.11
(b) Effects of frequency projection
SFI-STFT discriminator
As shown in Table 6(a), a shared SFI-STFT discriminator consistently outperforms separate rate-specific discriminators across all tasks. Because the SFI representation aligns TF structure across sampling rates, the unified discriminator receives more coherent supervision and produces more stable gradients. In contrast, separate discriminators see only partial bandwidth conditions, leading to weaker adversarial signals. These results confirm that the shared SFI design is more effective for multi-rate restoration.
Effects of frequency projection
Finally, we examine the influence of the frequency-projection (F-proj.) module. Introducing projection provides an explicit structural prior along the frequency axis, which stabilizes training and yields consistent improvements over the no-projection baseline across tasks. The difference between using shared and separate projections is relatively small, though the separate version exhibits less stable behavior in the super-resolution setting. More critically, the non-shared design is highly inefficient, expanding the model to 63.6M parameters. Given the similar performance and the large gap in model size, the shared frequency-projection module offers the most practical and efficient configuration.
Table 7: Ablation study on spectral loss. log1p denotes $\log(1+d)$ while s-log1p is the proposed scaled log1p $w\log(1+d/w)$ where $d$ is $\ell_{1}$ distance.
[…]
spectral loss PESQ MCD UTMOS PESQ MCD UTMOS
None 1.85 7.23 4.02 3.05 2.47 4.11
$\ell_{1}$-norm (mag.) 2.07 6.03 3.82 3.42 2.10 4.10
$\ell_{1}$-norm 2.23 5.70 3.76 3.48 1.86 4.07
$\ell_{2}$-norm 2.21 5.81 3.70 3.44 1.88 4.06
$\ell_{1}$-norm 2.19 5.89 3.71 3.35 2.23 3.91
log1p   ($w=1$) 2.25 5.72 3.79 3.53 1.83 4.06
s-log1p ($w=10^{-\hskip-1.42262pt3}$) 2.27 5.17 3.98 3.67 1.37 4.07
s-log1p ($w=10^{-\hskip-1.42262pt4}$) 2.01 5.94 4.07 3.40 2.43 4.10
s-log1p ($w=10^{-\hskip-1.42262pt3}$) 2.18 6.05 3.74 3.27 1.88 3.94
s-log1p (adap. $w_{tf}$) 2.29 4.96 4.10 3.70 1.29 4.10
Effects of scaled log-spectral loss
In Table 7, we assess whether auxiliary spectral losses provide benefits. Using perceptual loss alone leads to less stable optimization, whereas adding any spectral term consistently improves performance, confirming the importance of spectral constraints. Among regression-based losses, the $\ell_{1}$ loss on complex STFT components (magnitude, real, imaginary) outperforms magnitude-only variants by better preserving signal fidelity. Replacing $\ell_{1}$ with $\ell_{2}$ slightly degrades performance, likely due to oversmoothing. These results indicate that perceptual loss is essential for high-level quality but must be paired with an appropriate spectral objective.
We next compare log- and scaled log-spectral formulations. A plain log1p loss behaves similarly to $\ell_{1}$ because typical spectral distances are far below 1, keeping its gradient near 1. The proposed scaled log-spectral loss provides additional gains by adjusting gradient magnitude according to the target spectrum: suitable scale values balance well-aligned and poorly aligned regions, whereas overly small scales collapse gradients and damage performance. Removing perceptual loss noticeably harms both $\ell_{1}$ and s-log1p, and in this setting $\ell_{1}$ remains more stable, showing that s-log1p is not effective as a standalone objective. The best overall results arise when $w_{tf}$ is adaptively derived from the target magnitude, demonstrating that the proposed magnitude-adaptive scaling offers the most reliable trade-off between fine spectral detail and global coherence.
@@ 6 Conclusion
We presented TF-Restormer, a framework for speech restoration that explicitly addresses decoupled input–output sampling rates. By adopting a query-based asymmetric modeling strategy, the framework concentrates analysis on the observed input bandwidth while reconstructing missing high-frequency components through a lightweight decoder with frequency extension queries. This design removes reliance on conventional resampling and enables a single model to operate across arbitrary input–output rate pairs.
Extensive evaluations across diverse scenarios, including denoising, super-resolution, and compounded distortions, demonstrate balanced improvements in both signal fidelity and perceptual quality. Moreover, the shared SFI-STFT discriminator and the proposed scaled log-spectral loss ensure stable optimization under heterogeneous configurations. Overall, TF-Restormer establishes a unified framework for speech restoration under realistic and mismatched sampling-rate conditions.
Impact Statement
@@
Appendix A Details of Training Procedure
A.1 Simulation of training dataset
@@
A.2 Training Details for Unified Model
For pretraining, TF-Restormer was optimized with a batch size of 2 on a single NVIDIA RTX 6000 Ada 48GB GPU using AdamW (Loshchilov and Hutter, 2019). Pretraining was run for 200,000 steps on the VCTK dataset with 3-second utterances. Adversarial training was then applied for an additional 200,000 steps. We used a learning rate of 2.0e-4 with betas (0.9, 0.995), applying a decay of 0.9 every 10,000 steps after 100,000 steps during pretraining, and every 10,000 steps during adversarial training.
Both stages used a 5,000-step linear warm-up for the generator. In adversarial training, the discriminator was updated twice per generator step (without warm-up), using AdamW with betas (0.8, 0.999). Following multi-scale STFT discriminator designs (Défossez et al., 2023), we employed our multi-scale SFI-STFT discriminator with STFT window sizes of [20, 40, 60, 80, 100] ms to capture spectral details at multiple resolutions.
Across all ablation variants, validation loss plateaued around 60k–70k steps, and no architecture exhibited signs of overfitting. We observed that checkpoint selection within this plateau region led to negligible performance differences, indicating that the chosen training length is sufficient for convergence and provides a fair comparison across variants.
A.3 Training Details for dedicated model
VCTK+DEMAND
Since noise reduction does not require generating new speech components, prior work has shown that standard supervised learning is often sufficient. Therefore, in the fine-tuning stage we use small weights in adversarial loss ($\lambda_{\text{g}}=0.001$ and $\lambda_{\text{fm}}=0.01$ and the human-feedback perceptual loss ($\lambda_{\text{hf}}=10^{-5}$). The model is trained to perform pure denoising following the standard VCTK+DEMAND training partition. The input and output sampling rates are both fixed to 16 kHz, and thus no extension queries are used in this setting.
VCTK for Super-resolution
For super-resolution, the model is trained under the same protocol as conventional SR systems, using clean low-band inputs as supervision. Consequently, during adversarial fine-tuning the human-feedback loss is again applied with a small weight ($\lambda_{\text{hf}}=10^{-5}$), as the task primarily focuses on recovering missing high-frequency content.
Appendix B Details of Model Configuration
@@
B.1 Comparison of the model size and RTF
In Table 9, we compare model size and multiply-accumulate operations (MACs) for a 1-second-long input using ptflops package [^1] https://github.com/sovrasov/flops-counter.pytorch. We also measure real-time factor (RTF) measured on 4-second-long samples with an NVIDIA RTX 4090. Conventional models operate at fixed input–output sampling rates, which results in fixed MACs regardless of the task configuration. In contrast, TF-Restormer adapts its computation depending on the input and output rates $f_{E}$ and $f_{D}$.
@@ B.1 Comparison of the model size and RTF
Appendix C Simulation of VCTK Noisy Distorted Input
The noisy-distorted input from VCTK testset in Table 3 was generated by corrupting clean VCTK utterances with additive noise from DEMAND (Thiemann et al., 2013) and colored Gaussian noise, RIR samples from RWCP (Nakamura et al., 2000) and AIR (Jeub et al., 2009) for reverberation, and distortions such as clipping and band-limiting. Additional digital effects including audio codecs (MP3, OGG) were applied before resampling to various rates (8-48 kHz). This simulation aligns the training pipeline while maintaining samples partitioning of speech, noise, and RIR sources. As a result, the average SDR of input data (in case of $f_{E}=$16kHz) is 2.11dB from 2937 utterances. 998 utterances (about 34%) are below SDR=0 dB and 234 utterances (about 8%) are below SDR=-5 dB.
@@ Appendix C Simulation of VCTK Noisy Distorted Input
Appendix D Ablation Details
D.1 Effects of scaled log-spectral loss
@@
D.2 Illustration of encoder-decoder structure ablation
@@
Appendix E Extension Query Under Imbalanced Sampling-Rate Distributions
Table 11 analyzes whether extension-query tokens become undertrained when certain sampling rates appear too infrequently during training. For the input-rate ablation (top subtable), performance remains nearly identical to the balanced baseline as long as 8 kHz inputs constitute at least 10% of the data. The differences across {0.10, 0.25, 0.50} distributions are minimal in all target-rate settings, and even the best scores often occur at moderately imbalanced ratios. Noticeable degradation appears only in the extreme 1% case, where the model sees almost no examples of low-band inputs; this leads to modest but consistent drops, particularly for the largest gap ($8\to 44.1$ kHz).
A similar pattern is observed for the output-rate ablation (bottom subtable). When high-frequency target rates (44.1 kHz or 48 kHz) have extremely low probability (1–5%), reconstruction quality decreases for those specific targets, as seen in elevated LSD/MCD values. However, once each output rate is represented with a reasonable frequency (around 10% or more), the performance aligns closely with the uniformly balanced case, and the differences across distributions remain small.
Overall, these results show that extension-query undertraining affects performance only under highly skewed sampling-rate distributions. TF-Restormer remains robust as long as each rate appears with moderate frequency, and balanced or mildly imbalanced settings exhibit negligible differences from the baseline.
Table 11: Ablation study on input/output sampling-rate distributions to evaluate the robustness of extension-query training. Grey rows denote the default distribution used for baseline models.
$f_{E}$ (Hz) 8 $\to$ 16kHz 8 $\to$ 44.1kHz 16 $\to$ 48kHz
8k 16k LSD MCD NISQA LSD MCD NISQA LSD MCD NISQA
Input 3.36 11.38 1.91 3.64 11.47 1.91 3.48 11.37 1.73
0.01 0.99 1.48 5.83 4.01 1.46 7.12 4.41 1.15 2.82 4.55
0.10 0.90 1.21 2.82 4.48 1.24 3.20 4.46 1.17 2.85 4.56
0.25 0.75 1.16 2.78 4.49 1.18 3.08 4.52 1.18 2.86 4.54
0.50 0.50 1.14 2.74 4.50 1.17 3.05 4.52 1.20 2.89 4.53
(a) ablation of training $f_{E}$ distribution.
@@ Appendix E Extension Query Under Imbalanced Sampling-Rate Distributions
16k 24k 44.1k 48k LSD MCD NISQA LSD MCD NISQA LSD MCD NISQA LSD MCD NISQA
Input 3.36 11.38 1.91 3.49 11.60 1.91 3.64 11.47 1.91 3.48 11.37 1.73
0.49 0.49 0.01 0.01 1.15 2.73 4.51 1.17 3.05 4.52 1.36 4.01 4.49 1.51 6.83 4.01
0.45 0.45 0.05 0.05 1.15 2.74 4.51 1.16 3.05 4.51 1.20 3.16 4.49 1.27 3.00 4.46
0.40 0.40 0.10 0.10 1.16 2.76 4.50 1.16 3.04 4.51 1.21 3.15 4.49 1.23 2.94 4.49
0.25 0.25 0.25 0.25 1.16 2.78 4.49 1.18 3.08 4.52 1.18 3.08 4.52 1.18 2.86 4.54
(b) ablation of training $f_{D}$ distribution
Appendix F Evaluation of Dedicated Model on the URGENT Challenge
In the main paper, we reported non-intrusive MOS results on the URGENT blind test set using the unified TF-Restormer model. Since the blind set does not provide clean references, only MOS-based evaluation is possible. For completeness, we additionally train a dedicated URGENT model and report objective fidelity metrics on the non-blind validation set in the challenge.
For the dedicated URGENT configuration, we follow the official training recipe. Because the benchmark resamples all inputs to the target sampling rate regardless of their original rate, we remove high-frequency bins with negligible power prior to processing, which reduces redundant computation under this matched-rate setting. Apart from this preprocessing, the model is trained under the same conditions required by the challenge.
It is important to note that the URGENT benchmark emphasizes deterministic, fidelity-oriented enhancement (Sun et al., 2025; Chao et al., 2025; Rong et al., 2025). Intrusive metrics are heavily weighted, and top-ranked systems typically rely on large, deterministic architectures optimized exclusively for matched-rate denoising. In contrast, models that based on generative approach to improve perceptual quality generally achieve lower intrusive scores despite producing more natural listening quality.
Under such conditions, a dedicated TF-Restormer variant trained with the URGENT recipe shows moderate improvement in signal fidelity compared to the unified model; however, its performance remains below that of highly specialized deterministic and multi-stage systems and perceptual quality become significantly low. This outcome is expected: the strengths of TF-Restormer—arbitrary input–output sampling-rate handling, frequency-extension mechanisms, perceptually oriented losses, and adversarial training—do not align with the evaluation objectives of URGENT. Consequently, while the unified model yields strong non-intrusive MOS on the blind test set, the dedicated URGENT-trained version does not fully reflect the core advantages of our architecture.
Table 12: Evaluation results on URGENT 2025 (non-blind test set)
Model(Team) Model Rank intrusive signal fidelity Semantic fidelity Non-intrusive quality
Type PESQ ESTOI SDR MCD LSD sBERT SpkSim CAcc(%) UTMOS NISQA DNSMOS
Input - - 1.37 0.61 2.53 7.92 5.51 0.75 0.63 81.29 1.56 1.69 1.84
baseline D 9 2.43 0.80 11.29 3.32 2.84 0.86 0.80 84.96 2.11 2.89 2.94
Bobbsun D 1 2.95 0.86 14.33 3.01 2.83 0.91 0.85 88.92 2.09 3.22 2.88
USEM(rc) D 2 2.79 0.85 13.11 2.93 2.94 0.90 0.84 88.05 2.30 3.21 3.01
USEM-Flow(rc) G - 1.54 0.59 4.49 6.10 3.91 0.76 0.66 69.07 1.79 2.82 2.50
TS-URGENet(Xiaobin) D+G 3 2.74 0.84 13.06 3.30 3.08 0.89 0.84 87.94 2.16 3.24 2.92
alindborg D+G 10 1.99 0.76 7.49 4.51 3.73 0.84 0.77 81.70 2.49 3.96 3.28
wataru9871 G 13 1.36 0.56 -13.88 11.25 7.98 0.82 0.51 79.70 2.53 3.74 3.10
TF-Restormer (original) D+G - 1.71 0.71 4.19 4.84 4.32 0.84 0.73 80.21 3.57 4.51 3.25
TF-Restormer (dedicated) D+G - 2.60 0.83 11.78 3.18 2.91 0.88 0.82 84.28 2.47 3.43 3.06
Appendix G Comparison with Conventional Streaming Enhancement Models
We evaluate TF-Restormer-streaming against representative real-time denoising models on the VCTK+DEMAND test set, using PESQ, STOI, and the composite metrics CSIG, CBAK, and COVL commonly adopted in prior enhancement works. Unlike conventional streaming models, which are trained specifically for denoising under matched conditions, TF-Restormer-streaming inherits the full unified restoration and bandwidth-extension objective and is trained to handle reverberation, distortion, and bandwidth mismatch simultaneously. As a consequence, its model size and computational cost are substantially larger than lightweight denoisers such as NSNet2, DCCRN, or DeepFilterNet.
In terms of signal fidelity, specialized denoising models remain strong, with FRCRN and DeepFilterNet2 achieving the highest PESQ, CBAK, or STOI scores. Nevertheless, TF-Restormer-streaming attains competitive perceptual quality, achieving the highest CSIG score among all models and CBAK/COVL values close to the best discriminative systems. This is notable given that TF-Restormer-streaming is not optimized for denoising alone, but operates as a general-purpose restoration model that simultaneously handles reverberant, noisy, and bandwidth-limited inputs.
Overall, these results show that TF-Restormer-streaming is, to our knowledge, the first unified restoration and super-resolution model capable of streaming operation, while still providing signal fidelity comparable to denoising-oriented baselines. This demonstrates the feasibility of extending multi-rate, multi-distortion restoration models to real-time settings without sacrificing robustness.
Table 13: Comparison of TF-Restormer-streaming with existing real-time denoising models on the VCTK+DEMAND test set. We report standard enhancement metrics (PESQ, STOI, CSIG, CBAK, COVL) along with model size and MACs.
Model Size (M) MACs (G/s) PESQ CSIG CBAK COVL STOI
Noisy - - 1.97 3.34 2.44 2.63 0.921
NSNet2 (Braun et al., 2021) 6.2 0.43 2.47 3.23 2.99 2.90 0.903
DCCRN (Hu et al., 2020) 3.7 14.36 2.54 3.74 3.13 2.75 0.938
FullSubNet+ (Chen et al., 2022a) 8.7 30.06 2.88 3.86 3.42 3.57 0.940
FRCRN (Zhao et al., 2021) 10.3 12.3 3.21 4.23 3.64 3.73 -
DeepFilterNet (Schroter et al., 2022) 1.8 0.11 2.81 4.14 3.31 3.46 0.942
DeepFilterNet2 (Schröter et al., 2022) 2.3 0.356 3.08 4.30 3.40 3.70 0.941
TF-Restormer-streaming 19.0 214.5 2.89 4.37 3.41 3.68 0.937
Appendix H Evaluation on DNS Challenge Dataset
We further evaluate TF-Restormer on the 2020 DNS (Reddy et al., 2020) test sets to examine both signal fidelity and perceptual quality. Table 14 compares objective fidelity metrics against models specifically optimized for denoising with DNS training dataset. Since our unified model is trained to remove reverberation as well as noise, we report fidelity scores only on the “No Reverb” subset, where the clean references are aligned with our training objective. Under this setting, the unified TF-Restormer shows lower fidelity than DNS-targeted systems such as MFNet, USES, and TF-Locoformer. This gap is expected, as the unified model (i) is trained solely on VCTK, (ii) actively removes reverberation and other distortions, and (iii) incorporates perceptual objectives that may deviate from strict waveform fidelity. When trained in a DNS-specific manner without adversarial objectives, however, the dedicated TF-Restormer variant matches or surpasses prior systems, achieving competitive PESQ, STOI, and SI-SDR scores.
Table 15 presents perceptual DNSMOS scores on both “With Reverb” and “No Reverb” subsets. Here, discriminative models tend to preserve the input structure and thus achieve relatively conservative perceptual gains, as seen with Conv-TasNet and FRCRN. Generative approaches such as SELM, GenSE, MaskSR, and UniSE, which prioritize perceptual naturalness, obtain noticeably higher OVRL scores. TF-Restormer shows perceptual quality on par with these generative systems across all subsets, despite not being trained exclusively for perceptual enhancement. In both reverberant and non-reverberant conditions, it achieves strong SIG and BAK scores and matches the best OVRL scores among recent models, demonstrating that the proposed architecture can deliver high perceptual quality while maintaining reasonable signal fidelity.
Table 14: Comparison of TF-Restormer with previous models on 2020 DNS testsets in terms of signal fidelity. “No Reverb” subset is only compared as the proposed TF-Restormer is trained to remove reverberation.
System PESQ STOI(%) SI-SDR(dB)
Noisy 1.58 91.5 9.1
FullSubNet (Hao et al., 2021) 2.78 96.1 17.3
CTSNet (Li et al., 2021) 2.94 96.2 16.7
TaylorSENet (Li et al., 2022) 3.22 97.4 19.2
FRCRN (Zhao et al., 2021) 3.23 97.7 19.8
MFNet (Liu et al., 2023) 3.43 97.9 20.3
USES (Zhang et al., 2023) 3.46 98.1 21.2
TF-Locoformer (Saijo et al., 2024) 3.72 98.8 23.3
TF-Restormer 2.83 96.4 16.1
Table 15: Comparison of TF-Restormer with previous models on 2020 DNS testsets in terms of perceptual quality (DNS scores). “With Reverb” subset contains reverberation while “No Reverb” subset only involves noise. “D” and “G” denote discriminative and generative methods, respectively.
Model Type With Reverb No Reverb
SIG BAK OVRL SIG BAK OVRL
Noisy - 1.76 1.50 1.39 3.39 2.62 2.48
Conv-TasNet (Luo and Mesgarani, 2019) D 2.42 2.71 2.01 3.09 3.34 3.00
FRCRN (Zhao et al., 2021) D 2.93 2.92 2.28 3.58 4.13 3.34
SELM (Wang et al., 2024) G 3.16 3.58 2.70 3.51 4.10 3.26
MaskSR (Li et al., 2024) G 3.53 4.07 3.25 3.59 4.12 3.34
AnyEnhance (Zhang et al., 2025) G 3.50 4.04 3.20 3.64 4.18 3.42
GenSE (Yao et al., 2025) G 3.49 3.73 3.19 3.65 4.18 3.43
LLaSE-G1 (Kang et al., 2025) G 3.59 4.10 3.33 3.66 4.17 3.42
UniSE(Yan et al., 2025) G 3.67 4.10 3.40 3.67 4.14 3.43
TF-Restormer D+G 3.60 4.12 3.35 3.65 4.18 3.43
Appendix I Limitations and Future Works
Although TF-Restormer demonstrates balanced improvements in both signal fidelity and perceptual quality, several limitations remain.
First, when the input speech is extremely degraded, the uncertainty in the observed spectrum can lead to content or speaker artefacts. This reflects a fundamental trade-off in speech restoration: adversarial training promotes perceptually natural reconstructions by hallucinating plausible high-frequency details, whereas purely supervised objectives remain closer to the reference but often oversmooth, reducing naturalness.
Second, the model may exhibit mild rate-distribution sensitivity. As discussed in Appendix E, severely imbalanced exposure to certain sampling rates (e.g., <<1%) can lead to undertrained extension-query regions, particularly at the highest frequencies. Although moderate coverage ($\approx 10$%) is sufficient in practice, this highlights a remaining limitation of our arbitrary-rate formulation.
Third, because the model is trained on heavily distorted scenarios using limited single-language clean corpora, it may inherit language-specific biases, particularly due to the perceptual loss’s dependency on a pretrained speech model (e.g., WavLM). This could affect cross-lingual generalization to underrepresented phonetic patterns.
Finally, while TF-Restormer addresses universal restoration, it does not handle multi-speaker conditions such as speech separation or speaker extraction. The current formulation assumes a single target speaker and does not incorporate mechanisms for resolving overlapping speech.